# Running scalar spectral index in warm natural inflation

### v1

@@ Abstract
The validity of inflation models is mainly evaluated according to the consistency of the predicted scalar spectral index $n_{\mathrm{s}}$, the tensor scalar ratio $r$, and the running scalar spectral index $\alpha_{\mathrm{s}}$ with cosmic microwave background observations. In warm inflation (WI) scenarios, one can find exact analytical solutions for $\alpha_{\mathrm{s}}$ in principle, but long expressions may be obtained. Previous studies for WI scenarios have only shown approximate analytical solutions or numerical results for $\alpha_{\mathrm{s}}$. In this study, we present a general analytical expression of $\alpha_{\mathrm{s}}$ without approximation in WI. The obtained analytical expression of $\alpha_{\mathrm{s}}$ is used in the study of warm natural inflation (WNI). Although $n_{\mathrm{s}}$ and $r$ have been previously investigated, $\alpha_{\mathrm{s}}$ is omitted in previous studies on WNI. Our study of $\alpha_{\mathrm{s}}$ completes previous phenomenological studies on WNI. In particular, the lower limit of the symmetry-breaking scale in WNI becomes more concrete in this study.
@@ 1 Introduction
The inflationary paradigm [1, 2, 3, 4] is a cornerstone of standard cosmology, and several inflationary models consistent with observations have been developed [5]. In traditional cold inflation (CI) scenarios, the inflaton is assumed to have no interaction with other fields during inflation. In contrast, in warm inflation (WI) scenarios [6, 7, 8, 9], the inflaton interacts with other fields during inflation, and radiation production occurs. Because general relativity only requires that the inflaton energy density dominates for inflation to occur, substantial radiation energy densities present during inflation are allowed in principle. The validity of both CI and WI is mainly evaluated by the consistency of the predicted scalar spectral index $n_{\mathrm{s}}$ and the tensor scalar ratio $r$ with cosmic microwave background (CMB) observations.
In addition, the running scalar spectral index $\alpha_{\mathrm{s}}$ can be used to further check inflation models’ consistency. In CI scenarios, $\alpha_{\mathrm{s}}$ can be expressed in simple mathematical expressions using slow-roll parameters. In contrast, in WI scenarios, one can find exact analytical solutions for $\alpha_{\mathrm{s}}$ in principle but may get long expressions [10]. Therefore, previous studies on WI have only shown approximate analytical solutions [10, 11] or numerical results [12, 13] for $\alpha_{\mathrm{s}}$. Moreover, whereas $\alpha_{\mathrm{s}}$ has been studied in several CI models, it has been studied in only limited WI models, such as warm chaotic inflation, warm hilltop inflation, warm Higgs-like inflation, and minimal WI [12, 10, 13].
@@
  • The first objective is showing a general analytical expression of $\alpha_{\mathrm{s}}$ without approximation for two major dissipative coefficients: linear and cubic. By providing an analytical expression, even if it is mathematically redundant, we believe that $\alpha_{\mathrm{s}}$ will be studied across a broader range of WI models in the future.
@@ 2.2 Power spectrum
In the CI, the quantum fluctuation of the inflaton field is the only source of the primordial power spectrum. In the WI, the presence of dissipation affects the background dynamics of the inflaton field as well as its perturbations. Because of the temperature dependence of the dissipation coefficient, the inflaton and radiation fluctuations are coupled. The inflaton can be excited and obeys a Bose–Einstein distribution. When $T>H$, thermal fluctuations can dominate quantum density fluctuations. Consequently, the additional thermal effects significantly impact the primordial power spectrum. The dimensionless primordial scalar curvature power spectrum with the thermal effects is obtained as [43, 41, 40, 42]
@@
$n_{\mathrm{s}}-1$ $=4\frac{d\ln H}{dN}-2\frac{d\ln\dot{\phi}}{dN}+\left(1+2n_{\rm BE}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)^{-1}\left\{2n_{\rm BE}^{2}e^{H/T}\frac{H}{T}\left(\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right)\right.$
$\quad\left.+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\left[\frac{3+2\pi Q}{3+4\pi Q}\frac{d\ln Q}{dN}+\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right]\right\}+\frac{Q}{G(Q)}\frac{dG(Q)}{dQ}\frac{d\ln Q}{dN},$ (17)
@@ 2.3 Dissipation
$A$ $=\left(1+2n_{\rm BE}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)^{-1},$
$B$ $=n_{\rm BE}^{2}e^{H/T}\frac{H}{T}+\frac{6\sqrt{3}\pi Q(1+\pi Q)}{(3+4\pi Q)^{3/2}}\frac{T}{H},$ (31)
@@
$\quad\left.+\frac{4\sqrt{3}\pi Q}{(3+4\pi Q)^{3/2}}\frac{T}{H}\left[(21+9(3+2\pi)Q+26\pi Q^{2})\epsilon-2(1+Q)(6+5\pi Q)\eta\right.\right.$
$\quad\left.+(15+3(3+4\pi)Q+4\pi Q^{2})\sigma\right].$ (32)
@@
$\frac{d\ln A}{dN}$ $=-A\left\{2\cdot\frac{dn_{\rm BE}}{dN}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\left(\frac{3+2\pi Q}{3+4\pi Q}\frac{d\ln Q}{dN}+\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right)\right\},$ (33)
$\frac{dn_{\rm BE}}{dN}$ $=n_{\rm BE}^{2}e^{H/T}\frac{H}{T}\left(\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right),$ (34)
$\frac{d}{dN}\frac{dG}{dQ}$ $=Q\frac{d\ln Q}{dN}\frac{d^{2}G}{dQ^{2}},\quad\frac{d\ln G}{dN}=\frac{Q}{G}\frac{dG}{dQ}\frac{d\ln Q}{dN},$ (35)
@@ 3.2 Cubic dissipative coefficient
$n_{\mathrm{s}}-1$ $=-\frac{2(3+11Q)}{(1+Q)(1+7Q)}\epsilon+\frac{2}{1+7Q}\eta+\frac{16Q}{(1+Q)(1+7Q)}\sigma+\frac{2}{1+7Q}\frac{Q}{G}\frac{dG}{dQ}(5\epsilon-3\eta+4\sigma)$
$\quad+\frac{AD}{(1+Q)(1+7Q)}$ (48)
@@
$\quad+\frac{2(5\epsilon-3\eta+4\sigma)Q}{(1+7Q)G}\left\{\frac{d}{dN}\frac{dG}{dQ}+\frac{dG}{dQ}\left(\frac{1}{1+7Q}\frac{d\ln Q}{dN}-\frac{d\ln G}{dN}+\frac{5\epsilon_{,N}-3\eta_{,N}+4\sigma_{,N}}{5\epsilon-3\eta+4\sigma}\right)\right\}$
@@
$\quad+\frac{AD}{(1+Q)(1+7Q)}\frac{d\ln A}{dN}+\frac{AQ}{(1+Q)(1+7Q)}\left\{4n_{\rm BE}^{2}e^{H/T}\left[2(1+2Q)\epsilon-(1+Q)\eta+(1-Q)\sigma\right]\right.$
×$\qquad\times\left.\frac{1}{Q}\frac{H}{T}\left[2\cdot\frac{d\ln n_{\rm BE}}{dN}+\left(1+\frac{H}{T}\right)\left(\frac{d\ln H}{dN}-\frac{d\ln T}{dN}\right)\right]\right.$
$\quad-\left.\frac{1}{(1+Q)(1+7Q)}\frac{H}{T}\left[4n_{\rm BE}^{2}e^{H/T}\left(4(3+7Q+7Q^{2})\epsilon-7(1+Q)^{2}\eta\right.\right.\right.$
@@
$\quad+\frac{4\sqrt{3}\pi}{(3+4\pi Q)^{3/2}}\frac{T}{H}\left[Q\left((27+18\pi+52\pi Q)\epsilon-2(6+5\pi+10\pi Q)\eta+(9+12\pi+8\pi Q)\sigma\right)\right.$
$\qquad+\left.\frac{3-2\pi Q-3(7+16\pi)Q^{2}-70\pi Q^{3}}{(1+Q)(1+7Q)(3+4\pi Q)}\left((21+9(3+2\pi)Q+26\pi Q^{2})\epsilon\right.\right.$
$\qquad-\left.\left.\left.2(1+Q)(6+5\pi Q)\eta+(15+3(3+4\pi)Q+4\pi Q^{2})\sigma\right)\right]\frac{d\ln Q}{dN}\right\},$ (49)
respectively. As in the case of the linear dissipative coefficient, the relations in the CI are recovered in the limits $Q\rightarrow 0$ and $T\rightarrow 0$.
@@ 4.1 Slow-roll parameters and other key relations
[…]
The dimensionless dissipative ratio $Q$ can be numerically determined using the following relations:
@@ 4.2 Numerical analysis
The validity of WI (as well as CI) models can be evaluated according to the consistency of the predicted scalar spectral index $n_{\mathrm{s}}$, the tensor scalar ratio $r$, and the running scalar spectral index $\alpha_{\mathrm{s}}$ with CMB observations. For the $\Lambda$CDM model, the Planck 2018 TT,TE,EE+lowE+lensing data constrain the scalar spectral index and its running at 68% confidence level (CL) as follows [51]:
@@
[…]
@@ 5 Summary
$n_{\mathrm{s}}$ $=0.9587\pm 0.0056\ (0.9625\pm 0.0048),$ (63)
$\alpha_{\mathrm{s}}$ $=0.013\pm 0.012\ (0.002\pm 0.010),$ (64)
@@
at 68% CL for the $\Lambda$CDM model—the Planck 2018 TT(TT,TE,EE)+lowE+lensing data [39]. If these constraints are employed in our analysis, the numerical results in this paper will change. However, the running scalar spectral index $\alpha_{\mathrm{s}}$ must further be investigated in WI scenarios.

### v3

@@ Abstract
The validity of inflation models is mainly evaluated according to the consistency of the predicted scalar spectral index $n_{\mathrm{s}}$, the tensor scalar ratio $r$, and the running scalar spectral index $\alpha_{\mathrm{s}}$ with cosmic microwave background observations. In warm inflation (WI) scenarios, one can find exact analytical solutions for $\alpha_{\mathrm{s}}$ in principle, but long expressions may be obtained. Previous studies for WI scenarios have only shown approximate analytical solutions or numerical results for $\alpha_{\mathrm{s}}$. In this study, we present a general analytical expression of $\alpha_{\mathrm{s}}$ without approximation in WI. By providing an analytical expression, even if it is mathematically redundant, we believe that $\alpha_{\mathrm{s}}$ will be studied across a broader range of WI models in the future. The obtained analytical expression of $\alpha_{\mathrm{s}}$ is used in the study of warm natural inflation (WNI). Although $n_{\mathrm{s}}$ and $r$ have been previously investigated, $\alpha_{\mathrm{s}}$ is omitted in previous studies on WNI. Our study of $\alpha_{\mathrm{s}}$ completes previous phenomenological studies on WNI. In particular, the lower limit of the symmetry-breaking scale in WNI becomes more concrete in this study.
{history}
@@ 1 Introduction
The inflationary paradigm [1, 2, 3, 4] is a cornerstone of standard cosmology, and several inflationary models consistent with observations have been developed [5]. In traditional cold inflation (CI) scenarios, the inflaton is assumed to have no interaction with other fields during inflation. In contrast, in warm inflation (WI) scenarios [6, 7, 8, 9, 10], the inflaton interacts with other fields during inflation, and radiation production occurs. Because general relativity only requires that the inflaton energy density dominates for inflation to occur, substantial radiation energy densities present during inflation are allowed in principle. The validity of both CI and WI is mainly evaluated by the consistency of the predicted scalar spectral index $n_{\mathrm{s}}$ and the tensor scalar ratio $r$ with cosmic microwave background (CMB) observations.
In addition, the running scalar spectral index $\alpha_{\mathrm{s}}$ can be used to further check inflation models’ consistency. In CI scenarios, $\alpha_{\mathrm{s}}$ can be expressed in simple mathematical expressions using slow-roll parameters. In contrast, in WI scenarios, one can find exact analytical solutions for $\alpha_{\mathrm{s}}$ in principle but may get long expressions [11]. Therefore, previous studies on WI have only shown approximate analytical solutions [11, 12] or numerical results [13, 14, 15] for $\alpha_{\mathrm{s}}$. Moreover, whereas $\alpha_{\mathrm{s}}$ has been studied in several CI models, it has been studied in only limited WI models, such as warm chaotic inflation, warm hilltop inflation, warm Higgs-like inflation, and minimal WI [13, 11, 14].
@@
  • The first objective is showing a general analytical expression of $\alpha_{\mathrm{s}}$ without approximation for two major dissipative coefficients: linear and cubic. We also demonstrate that the analytical expressions of $\alpha_{\mathrm{s}}$ is consistent with the results from numerical calculation in Ref. [15]. By providing an analytical expression, even if it is mathematically redundant, we believe that $\alpha_{\mathrm{s}}$ will be studied across a broader range of WI models in the future.
@@ 2.2 Power spectrum
In the CI, the quantum fluctuation of the inflaton field is the only source of the primordial power spectrum. In the WI, the presence of dissipation affects the background dynamics of the inflaton field as well as its perturbations. Because of the temperature dependence of the dissipation coefficient, the inflaton and radiation fluctuations are coupled. The inflaton can be excited and obeys a Bose–Einstein distribution. When $T>H$, thermal fluctuations can dominate quantum density fluctuations. Consequently, the additional thermal effects significantly impact the primordial power spectrum. The dimensionless primordial scalar curvature power spectrum with the thermal effects is obtained as [43, 44, 42, 45, 46, 47]
@@
$n_{\mathrm{s}}-1$ $=4\frac{d\ln H}{dN}-2\frac{d\ln\dot{\phi}}{dN}+\left(1+2n_{\rm BE}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)^{-1}$
$\times\left\{2n_{\rm BE}^{2}e^{H/T}\frac{H}{T}\left(\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right)\right$.
$+\left.\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\left[\frac{3+2\pi Q}{3+4\pi Q}\frac{d\ln Q}{dN}+\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right]\right\}+\frac{Q}{G(Q)}\frac{dG(Q)}{dQ}\frac{d\ln Q}{dN},$ (17)
@@ 2.3 Dissipation
$A=\left(1+2n_{\rm BE}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)^{-1},\quad B=n_{\rm BE}^{2}e^{H/T}\frac{H}{T}+\frac{6\sqrt{3}\pi Q(1+\pi Q)}{(3+4\pi Q)^{3/2}}\frac{T}{H},$ (31)
@@
$+\frac{4\sqrt{3}\pi Q}{(3+4\pi Q)^{3/2}}\frac{T}{H}\left[(21+9(3+2\pi)Q+26\pi Q^{2})\epsilon\right$.
$\quad\left.-2(1+Q)(6+5\pi Q)\eta+(15+3(3+4\pi)Q+4\pi Q^{2})\sigma\right].$ (32)
@@
$\frac{d\ln A}{dN}=$ $-A\left\{2\cdot\frac{dn_{\rm BE}}{dN}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\left(\frac{3+2\pi Q}{3+4\pi Q}\frac{d\ln Q}{dN}+\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right)\right\},$
$\frac{dn_{\rm BE}}{dN}=$ $n_{\rm BE}^{2}e^{H/T}\frac{H}{T}\left(\frac{d\ln T}{dN}-\frac{d\ln H}{dN}\right),\quad\frac{d}{dN}\frac{dG}{dQ}=Q\frac{d\ln Q}{dN}\frac{d^{2}G}{dQ^{2}},$
$\quad\frac{d\ln G}{dN}=$ $\frac{Q}{G}\frac{dG}{dQ}\frac{d\ln Q}{dN},$ (33)
@@ 3.2 Cubic dissipative coefficient
$n_{\mathrm{s}}-1=$ $-\frac{2(3+11Q)}{(1+Q)(1+7Q)}\epsilon+\frac{2}{1+7Q}\eta+\frac{16Q}{(1+Q)(1+7Q)}\sigma$
$+\frac{2}{1+7Q}\frac{Q}{G}\frac{dG}{dQ}(5\epsilon-3\eta+4\sigma)+\frac{AD}{(1+Q)(1+7Q)}$ (46)
@@
[…]
$\quad\left.\left.+\frac{5\epsilon_{,N}-3\eta_{,N}+4\sigma_{,N}}{5\epsilon-3\eta+4\sigma}\right)\right\}$
@@
[…]
$\qquad\left.-(1+Q)\eta+(1-Q)\sigma\right]\times\left.\frac{1}{Q}\frac{H}{T}\left[2\cdot\frac{d\ln n_{\rm BE}}{dN}+\left(1+\frac{H}{T}\right)\left(\frac{d\ln H}{dN}-\frac{d\ln T}{dN}\right)\right]\right$.
[…]
@@
[…]
$\quad\left.+(9+12\pi+8\pi Q)\sigma\right)$
[…]
$\quad-\left.\left.\left.2(1+Q)(6+5\pi Q)\eta+(15+3(3+4\pi)Q+4\pi Q^{2})\sigma\right)\right]\frac{d\ln Q}{dN}\right\},$ (47)
respectively. As same as the case of the linear dissipative coefficient, the relations in the CI are recovered in the limits $Q\rightarrow 0$ and $T\rightarrow 0$.
3.3 Comparison with a previous study
We demonstrate that the analytical expression of $\alpha_{s}$ obtained in this study is consistent with the results from numerical calculation by WI2easy [15]. Figure 1 shows the predictions of $\alpha_{s}$ in the warm chaotic inflation potential $V=\lambda\phi^{4}/4$. The left (right) panel shows $\alpha_{s}$ vs $Q_{*}$, dimensionless dissipative ratio at the Hubble radius crossing, for the linear (cubic) dissipative coefficient. In the both panels, the blue curve depicts the results from WI2easy and the orange curve shows the prediction using by the analytical expressions in this study. Since we use approximate expressions of $G(Q)$ in Eqs. (39) or (45) for linear or cubic dissipative coefficient but WI2easy always use the numerically derived data for $G(Q)$, a little differences are observed in the blue and orange curves. We can conclude that the analytical expression of $\alpha_{s}$ is consistent with the WI2easy. This observation shows the availability of not only the analytical expression of $\alpha_{s}$ but also the WI2easy code.
Figure 1: Comparison with a previous study.
@@ 4.1 Slow-roll parameters and other key relations
$N=\int_{\tilde{\phi}_{*}}^{\tilde{\phi}_{\rm end}}(1+Q)\tilde{f}\frac{1+\cos(\tilde{\phi}/\tilde{f})}{\sin(\tilde{\phi}/\tilde{f})}d\tilde{\phi}$. (53)
The dimensionless dissipative ratio $Q$ can be determined using the following relations  [^3] Eq.(3.8) in [20] may be corrected as Eqs. (54) and (56).:
@@ 4.2 Numerical analysis
[…]
@@
Before presenting the results of the numerical analysis for WNI, we present the CNI results ($Q=0$ case) for $\tilde{f}=\{4,5,7,50\}$ and N=40N=406060. In the left panel in Fig. 2, the light gray (light green) and gray (green) regions show the constraints of $n_{\mathrm{s}}$ and $r$ from Planck 2018 (Planck with BICEP/Keck) at 98% and 68% CL. The colored curves denote the prediction of $n_{\mathrm{s}}$ and $r$. The right panel shows the prediction of $n_{\mathrm{s}}$ and $\alpha_{\mathrm{s}}$. The horizontal green band shows the constraint of $\alpha_{\mathrm{s}}$ from Planck 2018 at 68% CL. The two vertical lines denote the region of $n_{\mathrm{s}}$ from Plank 2018. From the left panel, the traditional CNI model may almost be excluded with observation. Therefore, a nonzero dissipation effect $Q\neq 0$ may be required to save the idea of the natural inflation models and consider WNI to be significant.
@@ 5 Summary
$n_{\mathrm{s}}$ $=0.9587\pm 0.0056\ (0.9625\pm 0.0048),$
$\alpha_{\mathrm{s}}$ $=0.013\pm 0.012\ (0.002\pm 0.010),$
@@
[…]
Acknowledgement
The author, T. K., thanks Rudnei O. Ramos for useful discussions.